### v2

@@ Abstract
The backpropagation algorithm remains the dominant and most successful method for training deep neural networks (DNNs). At the same time, training DNNs at scale comes at a significant computational cost and therefore a high carbon footprint. Converging evidence suggests that input decorrelation may speed up deep learning. However, to date, this has not yet translated into substantial improvements in training efficiency in large-scale DNNs. This is mainly caused by the challenge of enforcing fast and stable network-wide decorrelation. Here, we show for the first time that much more efficient training of very deep neural networks using decorrelated backpropagation is feasible. To achieve this goal we made use of a novel algorithm which induces network-wide input decorrelation using minimal computational overhead. By combining this algorithm with careful optimizations, we achieve a more than two-fold speed-up and higher test accuracy compared to backpropagation when training a 18-layer deep residual network. This demonstrates that decorrelation provides exciting prospects for efficient deep learning at scale.
keywords:
Deep learning, Backpropagation, Efficient AI, Decorrelation, Whitening
@@ 1 Introduction
Modern AI relies heavily on deep learning (DL), which refers to the training of very deep neural network (DNN) models using massive datasets deployed on high-performance compute clusters [1]. The established way of implementing learning in artificial neural networks is through the backpropagation (BP) algorithm [2, 3]. BP is a gradient-based method that implements reverse-mode automatic differentiation to compute the gradients needed for parameter updating in neural networks [4].
@@
At the same time, training of DNNs consisting of many layers of nonlinear transformations is computationally expensive and has been associated with significant energy consumption at a global scale [5, 6, 7, 8, 9, 10]. Hence, we are in need of green AI solutions which can significantly reduce the energy consumption of modern AI systems [11, 12, 13]. Patterson et al. [14] describe several ways to reduce the carbon consumption of deep learning, focusing mostly on more effective deployment methods. Another key reason for deep learning’s large carbon consumption, however, is the inefficiency of the backpropagation algorithm itself. BP requires many iterations of gradient descent steps until parameters converge to their optimal values. For example, training a large GPT model can easily take several weeks on a large compute cluster [15].
@@
In this paper, we show that deep learning can be made much more efficient by enforcing decorrelated inputs throughout the network. Decorrelation has previously been proposed to make credit assignment more efficient in neural networks [16]. Intuitively, if a network layer’s inputs have highly correlated features, it will be more difficult for the learning algorithm to perform credit assignment as it is now unclear if a change in the loss should be attributed to feature $i$ or feature $j$ in case both are correlated. Previous work has indeed shown that promoting decorrelation positively impacts training in relatively shallow networks [17].
Similar results have been obtained when inputs are whitened such that inputs are not only decorrelated but also forced to have unit variance [18]. In [19], whitening has been shown to have a positive impact in training deep networks, resulting in faster convergence when measured by number of epochs and somewhat lower test error compared to regular BP. However, due to computational overhead this did not yet translate into a substantial reduction in wall-clock training time. Desjardins et al. [20] also demonstrate that whitening has benefits when training DNNs, though this did not result in faster convergence to the minimal validation loss compared to standard BP using batch normalisation. At the same time, [21] suggest that strict whitening can have a negative impact on generalization performance.
Hence, while the potential benefits of decorrelation and whitening are encouraging, their application for significantly reducing real-world training times in modern deep neural networks remains to be shown. This has two main reasons. First, enforcing network decorrelation during training is associated with significant computational overhead, thereby eliminating any gains in training efficiency. Second, it is not obvious how to effectively implement decorrelation in a stable manner in deep (convolutional) neural networks.
@@
In this paper we demonstrate, for the first time, that training of large-scale DNNs can be made much more efficient through the development of a novel decorrelated backpropagation (DBP) algorithm. DBP combines automatic differentiation with an efficient iterative local learning rule which effectively decorrelates layer inputs across the network. This iterative decorrelation procedure was previously introduced by Ahmad et al. [17] and has been shown to speed up learning in limited-depth fully-connected neural networks. Here, we extend the procedure to more efficiently and effectively train deep residual networks [22]. This is achieved by making the procedure suitable for application in convolutional layers and ensuring that decorrelation remains stable across layers. Furthermore, we show how the same algorithm can be used to whiten layer inputs, allowing us to compare the efficacy of decorrelation and whitening in DNNs.
In the following, we show that DBP yields a two-fold reduction in training time, while achieving better performance compared to regular BP. Hence, widespread application of our approach can yield a substantial reduction in the carbon consumption of modern deep learning.
@@ 2.1 Decorrelated backpropagation
$\bm{\theta}\leftarrow\bm{\theta}-\eta\nabla_{\bm{\theta}}\mathcal{L}\,,$ (1)
@@
Decorrelated backpropagation (DBP) proceeds as regular backpropagation but additionally enforces decorrelated inputs to layers in the network. Since decorrelation is imposed independently in each layer, we concentrate on describing the decorrelation procedure for one such layer, implementing a nonlinear transformation
@@
with $\mathbf{W}\in\mathbb{R}^{d\times d}$ the weights and $\mathbf{x}\in\mathbb{R}^{d}$ the input. This transformation may be either the kernel function of a convolutional layer applied to a (flattened) input patch or the transformation in a fully-connected layer. Note further that we ignore biases without loss of generality since these can be represented using weights with fixed constant input. Input decorrelation refers to the property that $\langle\mathbf{x}\mathbf{x}^{\top}\rangle$ is diagonal for inputs $\mathbf{x}$, where the expectation is taken over the data distribution. In case of whitened inputs, $\langle\mathbf{x}\mathbf{x}^{\top}\rangle$ is assumed to be the identity matrix [23]. We will refer to the former as the covariance constraint and the latter as the variance constraint.
To ensure that the input $\mathbf{x}$ is decorrelated, we assume that it is the result of a linear transform
$\mathbf{x}=\mathbf{R}\mathbf{z}\,,$ (3)
where the decorrelating matrix $\mathbf{R}$ transforms the correlated input $\mathbf{z}$ into a decorrelated input $\mathbf{x}$. Hence, the transformation in a layer is described by
@@ 2.2 Decorrelation learning rule
To implement decorrelated backpropagation, we need to update individual decorrelation matrices for each of the layers in the network. This can be achieved by minimizing $K$ decorrelation loss functions in parallel to minimizing the BP loss with the aim of accelerating learning speed.
One way to achieve whitening as a restricted form of decorrelation is via established approaches such as zero-phase component analysis (ZCA) [24, 25]. However, naive application of ZCA is prohibitively costly in practice since it requires application of the algorithm to the inputs of each layer for all of the examples in the dataset at each step of gradient descent. More efficient batch implementations of whitening have been shown to have a positive impact on neural network training [18, 20, 19] but this did not yet yield a substantial reduction in training time in large-scale models when considering the moment at which minimal validation loss is achieved. See Appendix A for a complexity analysis of the different approaches.
To enable efficient decorrelation and whitening, we extend a recently introduced approach that enables more efficient network-wide decorrelation [17] and make this suitable for application in large-scale models. To derive our decorrelation learning rule, we start by defining the total decorrelation loss $\langle L\rangle$ as the average decorrelation loss across layers. The layer-specific decorrelation loss is defined as $L=\langle\ell(\mathbf{x})\rangle$, where the expectation is taken over the empirical distribution and $\ell(\mathbf{x})=\sum_{i=1}^{d}\ell_{i}(\mathbf{x})$ is the sum over unit-wise losses $\ell_{i}$. We define the unit-wise loss as
$\ell_{i}(\mathbf{x})=(1-\kappa)\frac{1}{2}\sum_{j\neq i}(x_{i}x_{j})^{2}+\kappa\frac{1}{4}(x_{i}^{2}-1)^{2}\,,$ (5)
where the regularization parameter $\kappa$ interpolates between the covariance constraint ($\kappa=0$) and the unit variance constraint ($\kappa=1$) such that that $\kappa\in(0,1)$ promotes whitening. It follows that we may write the layer-wise decorrelation loss compactly as
$\ell(x)=\left\langle(1-\kappa)\textrm{Tr}(\mathbf{C}\mathbf{C}^{\top})+\kappa\textrm{Tr}(\mathbf{V}\mathbf{V}^{\top})\right\rangle\,,$ (6)
@@
$\frac{\partial\ell_{i}}{\partial x_{i}}=(1-\kappa)\sum_{j\colon j\neq i}\left(x_{i}x_{j}\right)x_{j}+\kappa(x_{i}^{2}-1)x_{i}\,.$ (8)
If we vectorize across units, we obtain $\nabla_{\mathbf{x}}\ell=(1-\kappa)\mathbf{C}\mathbf{x}+\kappa\mathbf{V}\mathbf{x}$. If we plug this into the update step in Eq. 7, we obtain
$\mathbf{x}\leftarrow\mathbf{x}-\epsilon\left((1-\kappa)\mathbf{C}+\kappa\mathbf{V}\right)\mathbf{x}\,.$ (9)
Ultimately, we aim to derive an update rule for $\mathbf{R}$ instead of $\mathbf{x}$. This can be obtained using the identity $\mathbf{x}=\mathbf{R}\mathbf{z}$ since this allows us to write $\mathbf{R}\mathbf{z}\leftarrow\mathbf{R}\mathbf{z}-\epsilon\left((1-\kappa)\mathbf{C}+\kappa\mathbf{V}\right)\mathbf{R}\mathbf{z}$. Dividing by $\mathbf{z}$ on both sides and averaging over input samples we obtain our decorrelation learning rule
@@
This decorrelation rule allows for efficient batch updating of $\mathbf{R}$ since it only requires computing the decorrelated input covariance followed by multiplication with the decorrelation matrix. Note that for $\kappa=0$ we only decorrelate the input and for $\kappa=1$ we only enforce the unit variance constraint. For $10<\kappa<$ 1 we include both constraints to induce input whitening. Figure 1 shows effective decorrelation and whitening on two-dimensional input data.
@@
2.3 Application to deep residual networks
@@
Effective training of deep neural networks using DBP requires several modifications that improve algorithm stability, learning speed and allow application to convolutional rather than fully connected layers. As a canonical example, we consider an 18-layer deep residual network (ResNet). A ResNet consists of multiple residual blocks that implement modern network components such as convolutional layers and residual connections [22]. In the following, we describe the modifications that are required for effectively applying DBP in deep neural networks.
@@
Figure 2: Implementation of decorrelated backpropagation in residual networks. a) Residual blocks as implemented by our networks. b) Patch-wise flattening of the inputs with a flattened dimension of $d=M\times M\times C_{in}$. c) Decorrelating/whitening transform of the data by decorrelation matrix $\mathbf{R}$. d) 1 $\times$ 1 convolution operation with the weights $\mathbf{W}$ on the decorrelated input patches.
@@
First, we apply the decorrelation learning rule to convolutional layers as follows. Consider a convolutional layer with input size $S=M_{in}\times M_{in}\times C_{in}$, where $M_{in}\times M_{in}$ is the size of the feature map and $C_{in}$ is the channel dimension. Naive implementation of decorrelation would require updating a $S\times S$ decorrelation matrix $\mathbf{R}$, which is computationally too prohibitive in case of large feature maps. Instead, as shown in Fig. 2, we do not decorrelate the layer’s entire input, but only the local image patches, so that for an image patch with dimensionality $D=K\times K\times C_{in}$, matrix $\mathbf{R}$ would be of much smaller size $D\times D$. The output of this patchwise decorrelation operation would then be $D$ for every image patch, after which we can apply a $1\times 1$ convolutional kernel of size $D\times C_{out}$ for the forward pass. This local decorrelation is sufficient to ensure more efficient learning of the kernel weights. Figure 2 depicts the structure of a residual block that is extended to implement decorrelating transforms.
@@
Second, applying the decorrelation learning rule can become costly when averaging over input samples in Eq. 10 in case of convolutional layers. This is due to the need to compute an outer product $\mathbf{X}\mathbf{X}^{\top}$ with $\mathbf{X}\in\mathbb{R}^{D\times p}$ the input to a layer, where $p$ is the product of the number of batch elements times the number of patches. We found empirically that updates of $\mathbf{R}$ need not use the entire mini-batch to learn the correlational structure of the data. Sampling just 10% of the samples in each batch yields almost identical performance while significantly reducing the computational overhead of computing $\mathbf{R}$’s updates.
@@
Third, another optimization that decreases the computational overhead of the decorrelating transform is to combine the matrices $\mathbf{W}$ and $\mathbf{R}$ into a condensed matrix $\mathbf{A}=\mathbf{W}\mathbf{R}$ prior to multiplying by the correlated input $\mathbf{z}$. The dimensionality of $\mathbf{z}$ is much larger than that of $\mathbf{W}$ due to the batch dimension, and thus we replace an expensive multiplication $\mathbf{W}(\mathbf{R}\mathbf{z})$ by a cheaper multiplication $(\mathbf{W}\mathbf{R})\mathbf{z}$, significantly reducing the time required for performing a forward pass. As an additional benefit, when training is complete, we only need to store the $\mathbf{A}$ matrices.

### v4

@@ Abstract
The backpropagation algorithm remains the dominant and most successful method for training deep neural networks (DNNs). At the same time, training DNNs at scale comes at a significant computational cost and therefore a high carbon footprint. Converging evidence suggests that input decorrelation may speed up deep learning. However, to date, this has not yet translated into substantial improvements in training efficiency in large-scale DNNs. This is mainly caused by the challenge of enforcing fast and stable network-wide decorrelation. Here, we show for the first time that much more efficient training of deep convolutional neural networks is feasible by embracing decorrelated backpropagation as a mechanism for learning. To achieve this goal we made use of a novel algorithm which induces network-wide input decorrelation using minimal computational overhead. By combining this algorithm with careful optimizations, we achieve a more than two-fold speed-up and higher test accuracy compared to backpropagation when training several deep networks up to a 50-layer ResNet model. This demonstrates that decorrelation provides exciting prospects for efficient deep learning at scale.
@@ 1 Introduction
Modern AI relies heavily on deep learning (DL), which refers to the training of very deep neural network (DNN) models using massive datasets deployed on high-performance compute clusters (LeCun et al., 2015). The established way of implementing learning in artificial neural networks is through the backpropagation (BP) algorithm (Linnainmaa, 1970; Werbos, 1974). BP is a gradient-based method that implements reverse-mode automatic differentiation to compute the gradients needed for parameter updating in neural networks (Baydin et al., 2018).
@@
At the same time, training of DNNs consisting of many layers of nonlinear transformations is computationally expensive and has been associated with significant energy consumption at a global scale (García-Martín et al., 2019; de Vries, 2023; Thompson et al., 2021; Luccioni et al., 2023b; Strubell et al., 2020; Luccioni et al., 2023a). Hence, we are in need of green AI solutions which can significantly reduce the energy consumption of modern AI systems (Gibney, 2022; Schwartz et al., 2020; Wang, 2021). Previous work describes several ways to reduce the carbon consumption of deep learning, focusing mostly on more effective deployment methods (Patterson et al., 2022). Another key reason for deep learning’s large carbon consumption, however, is the inefficiency of the backpropagation algorithm itself. BP requires many iterations of gradient descent steps until parameters converge to their optimal values. For example, training a large GPT model can easily take several weeks on a large compute cluster (Brown et al., 2020).
@@
In this paper, we show that the efficiency of deep learning can be substantially increased by enforcing that layer inputs remain decorrelated throughout the network. Decorrelation has previously been proposed to make credit assignment more efficient in neural networks (LeCun et al., 2012). Intuitively, if a network layer’s inputs have highly correlated features, it will be more difficult for the learning algorithm to perform credit assignment as it is now unclear if a change in the loss should be attributed to feature $i$ or feature $j$ in case both are correlated. Previous work has indeed shown that promoting decorrelation positively impacts training in relatively shallow networks (Ahmad et al., 2023). This advantage of decorrelation can be theoretically understood as a way to better align the gradient update with that of the natural gradient (Ahmad, 2024; Amari, 1998). Similar results have been obtained when inputs are whitened such that inputs are not only decorrelated but also forced to have unit variance (Luo, 2017). Whitening has also been shown to have a positive impact in training deep networks, resulting in faster convergence when measured by number of epochs (Huangi et al., 2018).
However, since decorrelation and whitening methods themselves incur computational overhead, convergence speedup as a function of the number of epochs is not reflected in reduction in wall-clock time or carbon consumption, which is the metric of choice in green AI. Furthermore, other work indicates that whitening may have a convergence speed that is similar to that of batch-normalized networks (Desjardins et al., 2015) and generalization performance may be negatively impacted (Wadia et al., 2021). Hence, while the theoretical benefits of decorrelation and whitening are encouraging, their practical application to significantly reduce real-world training time in modern deep neural networks remains to be shown.
@@
In this paper we demonstrate, for the first time, that training of DNNs can be substantially improved in practice by employing a novel decorrelated backpropagation (DBP) algorithm. DBP combines automatic differentiation with an efficient iterative local learning rule which effectively decorrelates layer inputs across the network. This iterative decorrelation procedure generalizes a method which was previously introduced in the context of biologically plausible learning, demonstrating faster convergence in limited-depth fully-connected neural networks (Ahmad et al., 2023). Here, we extend the method to allow efficiently and effectively training of deep convolutional neural networks (He et al., 2016).
Our method is compared on four DNN architectures trained on the ImageNet dataset. To ensure a fair comparison across methods, we include batch normalization and optimized hyper-parameter optimization for both methods separately.
We show that wall-clock time of DNN training can be significantly reduced, up to 50% of that of regular BP, also reducing carbon emissions. Furthermore, we show that test accuracy of DBP-trained networks is consistently higher than that of BP-trained networks. Overall, our results show that including a local decorrelation mechanism across network layers can substantially improve training of deep neural networks.
@@ 2.1 Decorrelated backpropagation
$\bm{\theta}\leftarrow\bm{\theta}-\eta\nabla_{\bm{\theta}}\mathcal{L}$ (1)
@@
Decorrelated backpropagation (DBP) operates as regular backpropagation but additionally enforces the input to network layers to be decorrelated. Since decorrelation is imposed independently in each layer, we concentrate on describing the decorrelation procedure for one such layer, implementing a nonlinear transformation
@@
with input $\mathbf{x}\in\mathbb{R}^{d}$, weights $\mathbf{W}\in\mathbb{R}^{m\times d}$ and output $\mathbf{y}\in\mathbb{R}^{m}$. The nonlinear transformation may be either the transformation in a fully-connected layer or the kernel function of a convolutional layer applied to a (flattened) input patch. We ignore biases without loss of generality since these can be represented using weights with fixed constant input.
Input decorrelation refers to the property that the second-moment matrix $\langle\mathbf{x}\mathbf{x}^{\top}\rangle$ is diagonal for inputs $\mathbf{x}$, where the expectation is taken over the data distribution. In case of whitened inputs, $\langle\mathbf{x}\mathbf{x}^{\top}\rangle$ is additionally assumed to be the identity matrix (Kessy et al., 2015). We will refer to the former as the covariance constraint and the latter as the variance constraint. To ensure that the input $\mathbf{x}$ is decorrelated, we assume that it is the result of a linear transform $\mathbf{x}=\mathbf{R}\mathbf{z}$, where the decorrelating matrix $\mathbf{R}$ transforms the correlated input $\mathbf{z}$ into a decorrelated input $\mathbf{x}$. Hence, the transformation in a layer is described by
@@ 2.2 Decorrelation learning rule
To implement decorrelated backpropagation, we need to update individual decorrelation matrices for each of the layers in the network. This can be achieved by minimizing $K$ decorrelation loss functions in parallel with minimizing the BP loss with the aim of accelerating learning speed. One way to achieve whitening as a restricted form of decorrelation is via established approaches such as zero-phase component analysis (ZCA) (Bell and Sejnowski, 1995, 1997). However, as described in Appendix A, naive application of ZCA is prohibitively costly in practice since it requires computing an expensive ZCA transform for the inputs of each layer at each gradient-descent step.
To enable efficient decorrelation and whitening in DNNs, we generalize a recently introduced approach that allows for efficient network-wide decorrelation (Ahmad et al., 2023).
To derive the decorrelation learning rule, we start by defining the total decorrelation loss as the average decorrelation loss across layers. The layer-specific decorrelation loss is defined as $L=\langle\ell(\mathbf{x})\rangle$, where the expectation is taken over the empirical distribution and $\ell(\mathbf{x})=\sum_{i=1}^{d}\ell_{i}(\mathbf{x})$ is the sum over unit-wise losses $\ell_{i}$. We define the unit-wise loss as
$\ell_{i}(\mathbf{x})=(1-\kappa)\frac{1}{2}\sum_{j\neq i}(x_{i}x_{j})^{2}+\kappa\frac{1}{4}(x_{i}^{2}-1)^{2}$ (4)
where the first term is the zero covariance constraint imposed by decorrelation and the second term is the unit variance constraint imposed by whitening. The whitening parameter $\kappa\in[0,1]$ trades off between the covariance constraint ($\kappa=0$) and the variance constraint ($\kappa=1$) and is whitening inducing when $\kappa>0$. we may now write the layer-wise decorrelation loss compactly as
$\ell(\mathbf{x})=\left\langle(1-\kappa)\textrm{Tr}(\mathbf{C}\mathbf{C}^{\top})+\kappa\textrm{Tr}(\mathbf{V}\mathbf{V}^{\top})\right\rangle$ (5)
@@
$\frac{\partial\ell_{i}}{\partial x_{i}}=(1-\kappa)\sum_{j\colon j\neq i}\left(x_{i}x_{j}\right)x_{j}+\kappa(x_{i}^{2}-1)x_{i}$ (7)
which, when vectorized across inputs, yields $\nabla_{\mathbf{x}}\ell=(1-\kappa)\mathbf{C}\mathbf{x}+\kappa\mathbf{V}\mathbf{x}$. If we now plug this into the update step in Eq. 6, we obtain $\mathbf{x}\leftarrow\mathbf{x}-\epsilon\left((1-\kappa)\mathbf{C}+\kappa\mathbf{V}\right)\mathbf{x}$. Note, however, that our goal is to derive an update rule for $\mathbf{R}$ rather than $\mathbf{x}$. This can be obtained using the identity $\mathbf{x}=\mathbf{R}\mathbf{z}$ since this allows us to write $\mathbf{R}\mathbf{z}\leftarrow\mathbf{R}\mathbf{z}-\epsilon\left((1-\kappa)\mathbf{C}+\kappa\mathbf{V}\right)\mathbf{R}\mathbf{z}$. Dividing by $\mathbf{z}$ on both sides and averaging over input samples we obtain the decorrelation learning rule
@@
This decorrelation rule allows for efficient batch updating of $\mathbf{R}$ since it only requires computing the decorrelated input covariance followed by multiplication with the decorrelation matrix.
Figure 1 illustrates effective decorrelation and whitening on two-dimensional input data using Eq. (8).
@@
2.3 Decorrelating deep convolutional neural networks
@@
Effective training of DNNs using DBP requires several modifications that improve algorithm stability, learning speed and allow application to convolutional rather than fully connected layers. As a canonical example, we consider a deep residual network (ResNet). A ResNet consists of multiple residual blocks that implement modern network components such as convolutional layers and residual connections (He et al., 2016). In the following, we describe the modifications that are required for effectively applying DBP in deep neural networks.
@@
Figure 2: Implementation of decorrelated backpropagation in residual networks. a) Residual blocks as implemented by our networks. b) Patch-wise flattening of the inputs with a flattened dimension of $d=M\times M\times C_{\text{in}}$. c) Decorrelating/whitening transform of the data by decorrelation matrix $\mathbf{R}$. d) 1 $\times$ 1 convolution operation with the weights $\mathbf{W}$ on the decorrelated input patches.
Decorrelating convolutional layers
@@
We apply the decorrelation learning rule to convolutional layers as follows. Consider a convolutional layer with input size $S=M_{\text{in}}\times M_{\text{in}}\times C_{\text{in}}$, where $M_{\text{in}}\times M_{\text{in}}$ is the size of the feature map and $C_{\text{in}}$ is the channel dimension. Naive implementation of decorrelation would require updating a $S\times S$ decorrelation matrix $\mathbf{R}$, which is computationally prohibitive in case of large feature maps. Instead, as shown in Fig. 2, we do not decorrelate the layer’s entire input, but only the local image patches, so that for an image patch with dimensionality $D=K\times K\times C_{\text{in}}$, matrix $\mathbf{R}$ would be of much smaller size $D\times D$. The output of this patchwise decorrelation operation would then be $D$ for every image patch, after which we can apply a $1\times 1$ convolutional kernel of size $D\times C_{\text{out}}$ for the forward pass. This local decorrelation is sufficient to ensure more efficient learning of the kernel weights. Figure 2 depicts the structure of a residual block that is extended to implement decorrelating transforms.
Subsampling during covariance estimation
@@
Second, applying the decorrelation learning rule can become costly when averaging over input samples in Eq. 8 in case of convolutional layers. This is due to the need to compute an outer product $\mathbf{X}\mathbf{X}^{\top}$ with $\mathbf{X}\in\mathbb{R}^{D\times p}$ the input to a layer, where $p$ is the product of the number of batch elements times the number of patches. We found empirically that updates of $\mathbf{R}$ need not use the entire mini-batch to learn the correlational structure of the data. Sampling just 10% of the samples in each batch yields almost identical performance while significantly reducing the computational overhead of computing $\mathbf{R}$’s updates, as shown in Appendix B.
Efficient matrix products
@@
A final optimization that decreases the computational overhead of the decorrelating transform is to combine the matrices $\mathbf{W}$ and $\mathbf{R}$ into a condensed matrix $\mathbf{A}=\mathbf{W}\mathbf{R}$ prior to multiplying by the correlated input $\mathbf{z}$. The dimensionality of $\mathbf{z}$ is much larger than that of $\mathbf{W}$ due to the batch dimension, and thus we replace an expensive multiplication $\mathbf{W}(\mathbf{R}\mathbf{z})$ by a cheaper multiplication $(\mathbf{W}\mathbf{R})\mathbf{z}$, significantly reducing the time required for performing a forward pass. As an additional benefit, when training is complete, we only need to store the $\mathbf{A}$ matrices.